# Towards Large-scale Generative Ranking

### v1

@@ Abstract
Generative recommendation has recently emerged as a promising paradigm in information retrieval. However, generative ranking systems are still understudied, particularly with respect to their effectiveness and feasibility in large-scale industrial settings. This paper investigates this topic at the ranking stage of Xiaohongshu’s Explore Feed, a recommender system that serves hundreds of millions of users. Specifically, we first examine how generative ranking outperforms current industrial recommenders. Through theoretical and empirical analyses, we find that the primary improvement in effectiveness stems from the generative architecture, rather than the training paradigm. To facilitate efficient deployment of generative ranking, we introduce RankGPT, a novel generative architecture for ranking. We validate the effectiveness and efficiency of our solution through online A/B experiments. The results show that RankGPT achieves significant improvements in user satisfaction with nearly equivalent computational resources compared to the existing production system.
@@ 1 Introduction
This paper investigates generative ranking systems in large-scale industrial scenarios. In particular, we first analyze potential sources of effectiveness in generative recommendations and then conduct experiments to validate our hypotheses using existing generative recommenders [23]. Experimental results demonstrate that the generative architecture is critical to achieving strong performance. However, current generative architectures tend to be inefficient, particularly in large-scale settings. To address this issue, we propose a novel architecture, RankGPT, to meet the requirements of large-scale training and inference. Online A/B experiments on the Explore Feed in Xiaohongshu [^1] Also known as RedNote: https://www.xiaohongshu.com/explore. (Figure 1 (left)), a recommender system serving hundreds of millions of users, demonstrate the effectiveness and efficiency of our proposed solution.
@@ 3.1 Key Mechanisms in Generative Paradigm
3.2 Comparison of Module Effects across Two Paradigms
@@ 4 Efficient Generative Ranking in Industrial Scenarios
The previous section highlights the importance of architecture in generative ranking. Not only is it crucial for performance, but it also impacts the overall design of future recommendation systems. This section introduces a novel generative architecture, RankGPT, to enable efficient training and inference on large-scale ranking tasks. RankGPT differs from existing work in two ways: item-action organization (Section 4.1) and position & time biases (Section 4.2).
Table 1 summarizes our empirical results in training performance. We use HSTU [23] as the baseline method. Transitioning to the action-oriented organization yields a speed-up of 78.7%, while adopting the proposed position & time biases yields a speed-up of 25.0%. In general, RankGPT achieves a total speed-up of 94.8% during training, with a slight improvement in AUC on the test set.
@@
+ All (RankGPT) +94.8% +0.0006
Table 1: Ablation study of RankGPT.
@@ 4.1 Item-Action Organization
Figure 2: Model architecture of RankGPT. Compared to existing approaches, e.g., HSTU [23], which adopt an item-oriented organization, our solution adopts an action-oriented organization.
@@ 4.2 Position & Time Biases
Figure 3: Illustration of input representation and candidate mask. (a) The input representation of RankGPT contains five kinds of embeddings. (b) Candidate Mask: We depict the mask structure for a user’s sequence with one historical behavior and two candidate items.
@@ 5.1 Generative Recommendation
Generative recommendation has emerged as a promising paradigm in information retrieval [13, 22, 20, 19, 1]. Unlike traditional recommendation approaches, generative recommendation aims to directly generate recommendations from a user’s historical behaviors by formulating the recommendation as a sequence generation task. TIGER [13] is the first generative retrieval framework. It first acquires hierarchical IDs of items by quantizing their semantic embeddings, and then trains a sequence-to-sequence model that predicts the semantic ID of the next item. ColaRec [20] and LETTER [18] study the problem of enhancing collaborative signals in quantization to integrate content knowledge and collaborative interaction. COBRA [22] addresses the information loss from the quantization by a coarse-to-fine generation mechanism, enabling more expressive generative modeling. Despite their advancements, the effectiveness and feasibility of generative ranking are still underexplored in large-scale scenarios in the real world. HSTU [23] is the first study to investigate the generative ranking task. It introduces an interleaved organization to predict the action by treating user actions as a new modality. In contrast, RankGPT views items as positional indicators and reformulates recommendation as an action-oriented generation problem. Furthermore, we systematically analyze the efficacy drivers in generative recommendation, yielding crucial insights for both understanding generative ranking paradigms and informing future architectural designs.
@@ 6 Online Experiments
To validate the effectiveness and feasibility of generative ranking in product scenarios, we conducted online experiments in Xiaohongshu’s Explore Feed. All models traced back more than three months of data and were trained in an online manner. For the control group, we randomly selected 10% of users of Xiaohongshu and applied the production ranking model. For the treatment group, we applied RankGPT to a randomly selected 10% of users. Each group contains tens of millions of users, and there is no overlap between groups.
@@
From the offline metrics, the improvements in both AUC and GAUC [29, 3] for primary tasks exceed 0.0020, while the improvements for other tasks range between 0.0005 and 0.0015. From the online metrics, we select four metrics to measure the online performance: time spent, the number of reads, the number of engagements, and lifetime over 7 days (LT7). The online A/B test result averaged over a 15-day experimental period is shown in Table 2, where RankGPT outperforms the production ranking in all metrics. Specifically, we observe that the improvement of RankGPT on cold-start items is particularly significant. We believe this improvement stems from RankGPT’s enhanced ability to leverage the world knowledge from content embeddings.
In terms of overhead, RankGPT and the production ranking model require a comparable amount of overall resources. Specifically, RankGPT incurs higher training costs but lower inference and storage costs. Furthermore, RankGPT demonstrates a significant improvement in P99 response time, outperforming the production ranking model by over 25%. This highlights the potential for further optimization in test-time scaling.
@@ 7 Conclusion
In this paper, we investigate the effectiveness and feasibility of generative ranking in large-scale industrial settings. Through both theoretical analysis and empirical results, we find that the generative architecture is the primary source of effectiveness in generative recommendation. We also introduce a novel generative architecture, named RankGPT, which treats items as positional information and focuses on iteratively predicting user behaviors to address the inefficiencies present in existing approaches. Extensive large-scale offline and online experiments demonstrate the effectiveness and efficiency of our proposed solution.

### v2

@@ Abstract
Generative recommendation has recently emerged as a promising paradigm in information retrieval. However, generative ranking systems are still understudied, particularly with respect to their effectiveness and feasibility in large-scale industrial settings. This paper investigates this topic at the ranking stage of Xiaohongshu’s Explore Feed, a recommender system that serves hundreds of millions of users. Specifically, we first examine how generative ranking outperforms current industrial recommenders. Through theoretical and empirical analyses, we find that the primary improvement in effectiveness stems from the generative architecture, rather than the training paradigm. To facilitate efficient deployment of generative ranking, we introduce GenRank, a novel generative architecture for ranking. We validate the effectiveness and efficiency of our solution through online A/B experiments. The results show that GenRank achieves significant improvements in user satisfaction with nearly equivalent computational resources compared to the existing production system.
@@ 1 Introduction
This paper investigates generative ranking systems in large-scale industrial scenarios. In particular, we first analyze potential sources of effectiveness in generative recommendations and then conduct experiments to validate our hypotheses using existing generative recommenders [23]. Experimental results demonstrate that the generative architecture is critical to achieving strong performance. However, current generative architectures tend to be inefficient, particularly in large-scale settings. To address this issue, we propose a novel architecture, GenRank, to meet the requirements of large-scale training and inference. Online A/B experiments on the Explore Feed in Xiaohongshu [^1] Also known as RedNote: https://www.xiaohongshu.com/explore. (Figure 1 (left)), a recommender system serving hundreds of millions of users, demonstrate the effectiveness and efficiency of our proposed solution.
@@ 3.1 Key Mechanisms in Generative Paradigm
3.2 Module Performance Comparison across Paradigms
@@ 4 Efficient Generative Ranking in Industrial Scenarios
The previous section highlights the importance of architecture in generative ranking. Not only is it crucial for performance, but it also impacts the overall design of future recommendation systems. This section introduces a novel generative architecture, GenRank, to enable efficient training and inference on large-scale ranking tasks. GenRank differs from existing work in two ways: item-action organization (Section 4.1) and position & time biases (Section 4.2).
Table 1 summarizes our empirical results in training performance. We use HSTU [23] as the baseline method. Transitioning to the action-oriented organization yields a speed-up of 78.7%, while adopting the proposed position & time biases yields a speed-up of 25.0%. In general, GenRank achieves a total speed-up of 94.8% during training, with a slight improvement in AUC on the test set.
@@
+ All (GenRank) +94.8% +0.0006
Table 1: Ablation study of GenRank.
@@ 4.1 Item-Action Organization
Figure 2: Model architecture of GenRank. Compared to existing approaches, e.g., HSTU [23], which adopt an item-oriented organization, our solution adopts an action-oriented organization.
@@ 4.2 Position & Time Biases
Figure 3: Illustration of input representation and candidate mask. (a) The input representation of GenRank contains five kinds of embeddings. (b) Candidate Mask: We depict the mask structure for a user’s sequence with one historical behavior and two candidate items.
@@ 5.1 Generative Recommendation
Generative recommendation has emerged as a promising paradigm in information retrieval [13, 22, 20, 19, 1]. Unlike traditional recommendation approaches, generative recommendation aims to directly generate recommendations from a user’s historical behaviors by formulating the recommendation as a sequence generation task. TIGER [13] is the first generative retrieval framework. It first acquires hierarchical IDs of items by quantizing their semantic embeddings, and then trains a sequence-to-sequence model that predicts the semantic ID of the next item. ColaRec [20] and LETTER [18] study the problem of enhancing collaborative signals in quantization to integrate content knowledge and collaborative interaction. COBRA [22] addresses the information loss from the quantization by a coarse-to-fine generation mechanism, enabling more expressive generative modeling. Despite their advancements, the effectiveness and feasibility of generative ranking are still underexplored in large-scale scenarios in the real world. HSTU [23] is the first study to investigate the generative ranking task. It introduces an interleaved organization to predict the action by treating user actions as a new modality. In contrast, GenRank views items as positional indicators and reformulates recommendation as an action-oriented generation problem. Furthermore, we systematically analyze the efficacy drivers in generative recommendation, yielding crucial insights for both understanding generative ranking paradigms and informing future architectural designs.
@@ 6 Online Experiments
To validate the effectiveness and feasibility of generative ranking in product scenarios, we conducted online experiments in Xiaohongshu’s Explore Feed. All models traced back more than three months of data and were trained in an online manner. For the control group, we randomly selected 10% of users of Xiaohongshu and applied the production ranking model. For the treatment group, we applied GenRank to a randomly selected 10% of users. Each group contains tens of millions of users, and there is no overlap between groups.
@@
From the offline metrics, the improvements in both AUC and GAUC [29, 3] for primary tasks exceed 0.0020, while the improvements for other tasks range between 0.0005 and 0.0015. From the online metrics, we select four metrics to measure the online performance: time spent, the number of reads, the number of engagements, and lifetime over 7 days (LT7). The online A/B test result averaged over a 15-day experimental period is shown in Table 2, where GenRank outperforms the production ranking in all metrics. Specifically, we observe that the improvement of GenRank on cold-start items is particularly significant. We believe this improvement stems from GenRank’s enhanced ability to leverage the world knowledge from content embeddings.
In terms of overhead, GenRank and the production ranking model require a comparable amount of overall resources. Specifically, GenRank incurs higher training costs but lower inference and storage costs. Furthermore, GenRank demonstrates a significant improvement in P99 response time, outperforming the production ranking model by over 25%. This highlights the potential for further optimization in test-time scaling.
@@ 7 Conclusion
In this paper, we investigate the effectiveness and feasibility of generative ranking in large-scale industrial settings. Through both theoretical analysis and empirical results, we find that the generative architecture is the primary source of effectiveness in generative recommendation. We also introduce a novel generative architecture, named GenRank, which treats items as positional information and focuses on iteratively predicting user behaviors to address the inefficiencies present in existing approaches. Extensive large-scale offline and online experiments demonstrate the effectiveness and efficiency of our proposed solution.